A baseball team has home games on Thursday and Saturday. The two games together earn $4226.00 for the team. Thursday's game generates $274.00 less than Saturday's game. How much money was taken in at each game?
How much money did Thursday's game generate?
How much money did Saturday's game generate?

Found 2 solutions by fractalier, stanbon:
Answer by fractalier(6550)

Call the money for each game, t and s.
From the facts we have
t + s = 4226
and
t = s - 274
Now substitute this into the first equation...we get
s - 274 + s = 4226
2s - 274 = 4226
2s = 4500
s = $2250
t is $274 less or
t = $1976

Answer by stanbon(75887)

A baseball team has home games on Thursday and Saturday. The two games together earn $4226.00 for the team. Thursday's game generates $274.00 less than Saturday's game. How much money was taken in at each game?
How much money did Thursday's game generate?
How much money did Saturday's game generate?
------
Equations:
t + s = 4226
t = s - 274
----------------------
Substitute for "t" and solve for "s":
s-274 + s = 4226
-------------
2s = 4500
s = $2250 (amt. taken in on Saturday)
----
t = 4226-2250 = $1976 (amt. taken in on Thursday)
------------
Cheers,
Stan H.
-------------